# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v4

@@ 9.9 How to enable general manipulation with dexterous capability?
Currently, there are two primary approaches to designing end-effector. The first approach customizes the end-effector for specific tasks. The second approach makes the multi-fingered end-effector resemble a human hand. The end-effector designed with the first approach is usually easier to control because it has fewer degrees of freedom compared to the end-effector designed with the second approach. In Billard and Kragic (2019), dexterity is divided into two types: extrinsic dexterity and intrinsic dexterity. Extrinsic dexterity involves using external support, such as friction, gravity, and contact surfaces, to compensate for the lack of degrees of freedom. Intrinsic dexterity refers to the hand’s ability to manipulate objects using its own degrees of freedom. Therefore, the first approach still has certain limitations for general manipulation. The design of the end-effector is related to robot proprioception $S_{robot}$ and this survey includes this topic in the State module.
@@
9.9.2 Model-Based Generation of a Sequence of Contact Points and Forces/Torques
The pipeline contact points and forces/torques mentioned above are derived sequentially. In practical applications, both can be optimized simultaneously (Xu et al. (2024)). However, fingertips do not always make contact with the object’s surface through optimization. This results in a non-smooth contact dynamic nature and creates challenges for planning in contact-rich manipulation (Pang (2023)). As shown in Fig. 11, $q^{a}$ and $\mu$ represent the actual and commanded positions of the actuated ball, while $q^{u}$ represents an unactuated box that slides along a rail with sufficient damping. Our goal is to push the box by the ball to transition from the current configuration $q^{u}$ to the goal configuration $q^{u}_{goal}$. Suppose $q^{u}_{+}$ is the steady-state position of the box after the ball is commanded to position $\mu$. We expect that $q^{u}_{+}$ and $q^{u}_{goal}$ are sufficiently close. Typically, $q^{u}_{+}=f(q^{u},q^{a},\mu)$ represents the system’s dynamic, which can be obtained through system identification (Suh et al. (2023)).
As shown in Fig. 11, a gradient exists only when the ball and the box are in contact, which presents challenges for optimization-based problem-solving methods. Current approaches for smoothing contact dynamics are mainly divided into analytic smoothing, first-order randomized smoothing, and zeroth-order randomized smoothing (Pang et al. (2023)). However, it is often necessary to consider frictional contact constraints, which can make the optimization problem non-convex. To address this issue, current methods primarily rely on Anitescu’s convex relaxation (Todorov et al. (2012); Pang et al. (2023); Todorov (2014); Anitescu (2006); Castro et al. (2022)). At the same time, the optimization should avoid robot joint limits, penetration between the hand and object, and self-penetration as much as possible (Pang and Tedrake (2021); Xu et al. (2024)). Although smoothed dynamics on trajectory optimization, such as iMPC (Suh et al. (2022) and LQR (Shirai et al. (2024)) are considered effective, these methods are only suitable when the goal is close to the initial configuration. When the goal is farther from the initial configuration, a more global strategy is required.
As shown in Fig. 11, there are two initial guesses: 1) the ball is on the left of the box, and 2) the ball is on the right of the box. Different initial guesses lead to different dynamic model. Choosing the wrong initial guess can prevent the optimization problem finding the optimal solution. To address this issue, the system dynamics need to be divided into contact modes, and a global strategy is required (Pang et al. (2023)). Fig. 11 includes four contact modes: left contact, left no contact, right contact, and right no contact. However, for a system with $n$ rigid bodies, there can be up to $C^{2}_{n}$ contact pairs. Each contact pair may have multiple contact modes, such as sticking, separation, and sliding, which can lead to an explosion of contact modes. Although Huang et al. (2021) considers kinematically feasible contact modes and reduces the number to $O(n^{d^{2}/2+2.5d})$(where d is the number of degrees of freedom and n is the number of contact points), the reduced number of contact modes is still too large to handle.
Figure 11: Non-smooth dynamic system.
@@ 9.9.2 Model-Based Generation of a Sequence of Contact Points and Forces/Torques
9.9.3 Learning-Based Generation of a Sequence of Contact Points and Forces/Torques
@@
One major challenge in data collection for dexterous manipulation lies in gathering data from multi-fingered end-effectors. Although model-based hard code method (Zhu et al. (2024)) can collect data on dexterous manipulation, they still require data analysis such as mutual information (Hejna et al. (2025)) and entropy (Zhu et al. (2024)) to assess the quality of the data. Additionally, for multi-scenario and multi-task data collection, teleoperation methods are less dependent on algorithm performance compared to model-based hard code methods. However online teleoperation requires a real-robot system, which is not portable and cannot achieve in-the-wild data collection. Therefore, current mainstream research focuses on directly tracking human hand motions during manipulation without controlling the real-robot.
Hand motion capture can be categorized into camera-based, IMU-based, and electromagnet-based methods. Using camera-based methods, human motion is estimated through pose estimation and retargeted to a multi-fingered end-effector. However, the accuracy of pose estimation needs improvement and occlusions can cause information loss. IMU-based methods are typically used to estimate wrist pose but miss fine-grained finger motion. Additionally, IMU used for pose estimation are prone to drifting over time. Electromagnet-based methods primarily use electromagnetic field gloves to estimate finger motion, but they lack information about wrist pose. At the same time, IMU-based methods and electromagnet-based methods both lack the vision data needed for policy training. DEXCAP (Wang et al. (2024b)) designs a portable hand motion capture system. It uses Rokoko gloves to capture finger motion. A Realsense T265 camera is added to the wrist of the gloves to monitor wrist pose. Additionally, a Realsense L515 LiDAR camera is mounted on the chest to sense the environment. However, the glove’s design lacks tactile feedback. Therefore, to collect data for a multi-fingered end-effector, hardware improvements are necessary.
@@
Two main learning-based methods for dexterous manipulation are imitation learning (Ze et al. (2024a)) and reinforcement learning (Ma et al. (2023b)). Imitation learning can use a visual encoder (in Sec. 6) for visuo-motor control. Diffusion policy (Chi et al. (2023)) adapts the concept of diffusion to visuo-motor control. It addresses challenges in visuo-motor control such as action multimodality, sequential correlation to accommodate high-dimensional action sequences. However, it remains uncertain whether diffusion policy is suitable for one-model to multiple tasks approach. It can also use an existing VLM for fine-tuning (in Sec. 7). Fine-tuning with a VLM allows a skill to work in an open world. This often performs better on unseen objects compared to visuo-motor control (Brohan et al. (2023)). Reinforcement learning offers exploration capability, which address suboptimal issues. This advantage distinguishes it from imitation learning. However, reinforcement learning is primarily trained in simulation. It still has limitations in addressing the sim-to-real challenge of complex tasks, such as pen-spinning. In Sec. 7, the use of foundation models to assist reinforcement learning is introduced. FAC (Ye et al. (2023a)) offers potential for training reinforcement learning in real-world environment, but it still lacks consideration of environment resets (Gupta et al. (2021)) and safety. Therefore, using foundation models to assist reinforcement learning in real-world training requires further exploration.
@@
Current learning methods each have their strengths and weaknesses (Zhang et al. (2024a)). Therefore, learning approaches for dexterous manipulation should integrate different methods. For example, diffusion policy can assist reinforcement learning in addressing high-dimensional action spaces issue, while reinforcement learning can help diffusion policy overcome issues with suboptimal and negative data. Additionally, the learning models should consider both inputs and outputs. The factors necessary for achieving dexterous manipulation are summarized in the ”Model-Based Generation of a Sequence of Contact Points and Forces/Torques”. ManiFoundation(Xu et al. (2024)) generates contact point and force heatmaps based on task goals, object material, object geometry and end-effector geometry. However, ManiFoundation lacks sensors such as tactile for real-time servoing of the manipulated object’s motion.
9.9.4 Whole-body Control
@@ 10 Conclusion
Then, we discuss the following points: (1) the logic and implementation strategies of the designed framework, (2) how to design the modules of post-conditions detection and post-hoc correction, due to limited research on foundation models for these areas, (3) the learning capability required for general manipulation, (4) the strengths and weaknesses of current RFMs, (5) how to use internet-scale video data for RFMs, (6) how to design low-cost hardware for collecting large-scale real-world manipulation data, (7) how to collect large-scale diversity and realism manipulation data in simulation, (8) comparative analysis of 2D and 3D-based methods, (9) how to make the general manipulation with the dexterous capability, and (10) how to establish a foundation models for manipulation benchmark.

### v5

@@ 9.6 How to use foundation models for End-effector Design?
Currently, there are two primary approaches to designing end-effector. The first approach customizes the end-effector for specific tasks. The second approach makes the multi-fingered end-effector resemble a human hand. The end-effector designed with the first approach is usually easier to control because it has fewer degrees of freedom compared to the end-effector designed with the second approach. In Billard and Kragic (2019), dexterity is divided into two types: extrinsic dexterity and intrinsic dexterity. Extrinsic dexterity involves using external support, such as friction, gravity, and contact surfaces, to compensate for the lack of degrees of freedom. Intrinsic dexterity refers to the hand’s ability to manipulate objects using its own degrees of freedom. Therefore, the first approach still has certain limitations for general manipulation.
@@
9.7 How to use foundation models for dexterous manipulation?
@@
One major challenge in data collection for dexterous manipulation lies in gathering data from multi-fingered end-effectors. Although model-based hard code method (Zhu et al. (2024)) can collect data on dexterous manipulation, they still require data analysis such as mutual information (Hejna et al. (2025)) and entropy (Zhu et al. (2024)) to assess the quality of the data. Additionally, for multi-scenario and multi-task data collection, teleoperation methods are less dependent on algorithm performance compared to model-based hard code methods. However online teleoperation requires a real-robot system, which is not portable and cannot achieve in-the-wild data collection. Therefore, current mainstream research focuses on directly tracking human hand motions during manipulation without controlling the real-robot (Wang et al. (2024b)).
@@
Two main learning-based methods for dexterous manipulation are imitation learning (Ze et al. (2024a)) and reinforcement learning (Ma et al. (2023b)). Imitation learning can use a visual encoder (in Sec. 6) for visuo-motor control. Diffusion policy (Chi et al. (2023)) adapts the concept of diffusion to visuo-motor control. It addresses challenges in visuo-motor control such as action multimodality, sequential correlation to accommodate high-dimensional action sequences. It can also use an existing RFMs for fine-tuning (in Sec. 7). Fine-tuning with a RFMs allows a skill to work in an open world. This often performs better on unseen objects compared to visuo-motor control (Brohan et al. (2023)).
Reinforcement learning offers exploration capability, which address suboptimal issues. This advantage distinguishes it from imitation learning. However, reinforcement learning is primarily trained in simulation. It still has limitations in addressing the sim-to-real challenge of complex tasks, such as pen-spinning. In Sec. 7, the use of foundation models to assist reinforcement learning is introduced. FAC (Ye et al. (2023a)) offers potential for training reinforcement learning in real-world environment, but it still lacks consideration of environment resets (Gupta et al. (2021)) and safety. Therefore, using foundation models to assist reinforcement learning in real-world training requires further exploration.
@@
Current learning methods each have their strengths and weaknesses (Zhang et al. (2024a)). Therefore, learning approaches for dexterous manipulation should integrate different methods. For example, diffusion policy can assist reinforcement learning in addressing high-dimensional action spaces issue, while reinforcement learning can help diffusion policy overcome issues with suboptimal and negative data. Additionally, the learning models should consider both inputs and outputs. The factors necessary for achieving dexterous manipulation are summarized in the Appendix.D.
9.8 How to use foundation models for whole-body control?
@@ 10 Conclusion
Then, we discuss the following points: (1) the logic and implementation strategies of the designed framework, (2) the learning capability required for general manipulation, (3) what foundation models bring for general manipulation, (4) how to use internet-scale video data for RFMs, (5) how to uses foundation models for post-conditions detection and post-hoc correction, (6) how to use foundation models for end-effector design, (7) how to use foundation models for dexterous manipulation, (8) how to use foundation models for whole-body control, and (9) how to establish a benchmark.